OpenStudy (anonymous):

The domain of the function f(x) is {0, -3, -6}. If f(x)= x+2/ 3-x what is the range? How do I solve this?

OpenStudy (anonymous):

f(x)=(x+2)/(x-3)?

OpenStudy (anonymous):

just do the substitution when x=0...y=-2/3 when x=-3...y=1/6 when x=-6...y=4/9 so the range is {-2/3,1/6,4/9}